OpenStudy (anonymous):

Determine whether the graph of y = x^2 + 4x − 7 has a maximum or minimum point, then find the maximum or minimum value. Minimum; (-11, -2) Maximum; (-11, -2) Maximum; (-2, -11) Minimum; (-2, -11)

OpenStudy (anonymous):

same as last one

OpenStudy (anonymous):

its different though

hartnn (hartnn):

simply taking derivative and equating to 0 gives x=-2,and since double derivative=2=+ve,its a minimum

OpenStudy (anonymous):

leading coefficient is 1, so it opens up and has a minimum but no max

OpenStudy (anonymous):

min is at the vertex, first coordinate of the vertex is always \[-\frac{b}{2a}\] in your case \(a=1,b=4\) and so \[-\frac{b}{2a}=-\frac{4}{2}=-2\]

OpenStudy (anonymous):

ok so its minimum the first one

OpenStudy (anonymous):

oh no no be careful

OpenStudy (anonymous):

\(-2\) is the first coordinate, not the second coordinate you find the second coordinate by replacing \(x\) by \(-2\) to see what you get

OpenStudy (anonymous):

their is only two minimuns

OpenStudy (anonymous):

ohhhh gotcha

OpenStudy (anonymous):

can u continue to help me please

OpenStudy (anonymous):

if \(x=-2\) then \[y=(-2)^2+4\times -2-7=-4-7=-11\] \[

OpenStudy (anonymous):

so the vertex is \((-2,-11)\)

OpenStudy (anonymous):

gotcha :)

OpenStudy (anonymous):

the question is worded incorrectly i don't know where it came from, but it is not worded right the MINIMUM value is a number, not a coordinate the minimum value is \(-11\)

OpenStudy (anonymous):

hmm idk